OpenStudy (anonymous):

What is the solution to the equation log2 x + log2 4 = 3 ?

OpenStudy (blockcolder):

Hint: \[\log_a x+\log_a y=\log_a (xy)\]

OpenStudy (anonymous):

I think it's 2?

OpenStudy (anonymous):

log 2 (4x) = 3 4x= 2^3 x=8/4 x=2